A card is randomly selected from an ordinary deck f playing cards.
a)what are the odds in favor of drawing a diamond ?
a) what are the odds against drawing a diamond ?
C) what are the odds in favor of getting a king ?
d) what are the odds against getting a king ?

Answer by htmentor(1343)

(a) Their are 52 cards of which 13 are diamonds.
Therefore P(diamond) = 13/52 = 1/4
(b) The odds against drawing a diamond are 1 - P(diamond) = 1 - 1/4 = 3/4
(c) There are 4 kings out of 52 cards.
So P(king) = 4/52 = 1/13
(d) Odds against getting a king are 1 - P(king) = 1 - 1/13 = 12/13